# Instructing Large Language Models for Low-Resource Languages: A Systematic Study for Basque

### v1

@@ Abstract
Instructing language models with user intent requires large instruction datasets, which are only available for a limited set of languages. In this paper, we explore alternatives to conventional instruction adaptation pipelines in low-resource scenarios. We assume a realistic scenario for low-resource languages, where only the following are available: corpora in the target language, existing open-weight multilingual base and instructed backbone LLMs, and synthetically generated instructions sampled from the instructed backbone. We present a comprehensive set of experiments for Basque that systematically study different combinations of these components evaluated on benchmarks and human preferences from participants. Our conclusions show that target language corpora are essential, with synthetic instructions yielding robust models, and, most importantly, that using as backbone an instruction-tuned model outperforms using a base non-instructed model, and improved results when scaling up. Using Llama 3.1 instruct 70B as backbone our model comes near frontier models of much larger sizes for Basque, without using any Basque data apart from the 1.2B word corpora. We release111Repository: https://github.com/hitz-zentroa/latxa-instruct code, models, instruction datasets, and human preferences to support full reproducibility in future research on low-resource language adaptation.
booktabs
@@ 1 Introduction
Specifically, this work systematically explores diverse strategies beyond the conventional pipeline for developing instruction-tuned models for low-resource languages, seeking to identify optimal adaptation paths for Basque as our primary case study (see Fig. 1). We deliberately constrain our exploration to resources either readily available or creatable using open models, avoiding reliance on distillation from commercial state-of-the-art systems. While our investigation focuses on a single language, our findings likely generalize to many similarly-resourced languages: Basque represents an ideal test case, ranking approximately 50th in Common Crawl with a presence approximately times smaller than English,222https://commoncrawl.github.io/cc-crawl-statistics/plots/languages.html and notably lacking pre-existing instruction datasets. This resource profile mirrors challenges faced by numerous other low-resource languages worldwide.
@@ 2 Related Work
Research on developing LLMs for under-resourced languages has explored various approaches, with varying degrees of success. Initial attempts to develop models from scratch for specific low-resource languages have proven challenging due to limited training data. Multilingual model development has emerged as a more promising strategy, with researchers leveraging cross-lingual transfer learning to improve performance Scao et al. (2023); Le Scao et al. (2022); Shliazhko et al. (2024). The most effective approach to date involves continued pre-training of existing multilingual models, which allows for language-specific adaptation while benefiting from the rich linguistic representations of larger training corpora Etxaniz et al. (2024b); Luukkonen et al. (2023); Tran et al. (2024). While progress has been made in developing these base models, optimal methods for instructing and fine-tuning them for under-resourced languages remain largely unexplored Gonzalez-Agirre et al. (2025); Martins et al. (2025); Üstün et al. (2024).
@@ 3 Resources
Instructing LLMs typically relies on two components: base (or foundational) LLMs and instruction datasets. For non-hegemonic languages, obtaining instruction datasets can be very challenging, particularly in low-resource language scenarios. In the case of Basque in particular, there are no manually generated, or even good quality automatically generated, large sets of instruction-answer pairs. Consequently, as shown in Fig. 1, our available resources are constrained to corpora on the target language, and base and instruct models for high-resource languages. From these limited resources, we derive the necessary components to create Basque instruction-tuned models through strategic combinations of synthetic data generation and model adaptation. In the following sections, we describe these seed resources and derivations.
@@ 3.1 Basque corpora
For the pre-training data, we have leveraged the corpora used to train Latxa, the first family of LLMs trained specifically for Basque Etxaniz et al. (2024b). This corpus comprises 4.3M of high-quality documents in Basque, roughly 3.5B Llama 3.1 tokens. Among the sources, it contains high-quality news data extracted using ad-hoc scrapers Artetxe et al. (2022), Wikipedia333The 20231101 dump corresponding to November 2023. and sources based on Common Crawl such as CulturaX Nguyen et al. (2024), Colossal OSCAR Abadji et al. (2022) and HLPT v1.1 de Gibert et al. (2024). This corpus comes normalized, deduplicated and filtered. The data is publicly available in the HuggingFace hub.444hf.co/datasets/HiTZ/latxa-corpus-v1.1 We will henceforth refer to this corpus as Corpus EU.
@@ 3.3 Instruction Sampling and Translation
To avoid external influences, we instead sample instructions directly from our InstructEN model. We generate the English instructions following Xu et al. (2025). Using this technique, we conditioned InstructEN to generate instructions of different types and tasks: general-purpose, code, math, arithmetic and translation. We generated a total of 4M English instructions, however, after a hyperparameter search, we found out that using just 1M instructions yielded better results overall. We share more details and examples of the process in Appendix A.
@@
We translated instructions sampled from InstructEN using few-shot prompting with Base EU. Existing machine translation systems for the English–Basque language pair (e.g., NLLB Costa-jussà et al. (2022)) are primarily trained on sentence-level textual data and often struggle with more complex inputs, including selectively translating natural language content embedded within code snippets. By leveraging an LLM like Base EU, which has been exposed to diverse data types, we obtained higher-quality translations for this setting. Moreover, using a model trained within our own experimental framework allows us to avoid introducing external factors into our pipeline. More details about the process and prompts used to translate the instructions are given in Appendix A.
@@ 4 Experimental Setup
We formalize our experimental setup as follows. Let be the set of backbone models and be the set of binary variables indicating whether to use Basque corpora, English instructions, and/or Basque instructions. The space of possible configurations is thus , where is the power set of , yielding theoretical combinations. Note that we explore training strategies that leverage both raw text and instruction data simultaneously. From the total of 24 combinations, we exclude redundant configurations where a model is retrained on data it was originally trained with. The resulting set of distinct instruction-tuned model variants therefore comprises 18 configurations: the original Llama 3.1 Instruct 8B (i.e., Instruct EN) and 17 new 8B-sized models. Table 3 in Appendix B provides a complete account of all model variants and their shorthand names. Additionally, we trained a 70B model following the configuration that performed best in preliminary benchmark evaluations.
@@ 5 Evaluation
We employed two complementary evaluation approaches to assess the impact of each instruction-tuning recipe. On the one hand, we used a selection of static benchmarks that evaluate specific model capabilities and knowledge through standardized tests. On the other hand, we conducted human evaluations through A/B testing (arena style) to capture qualitative aspects of model performance.
@@ 5.1 Static Benchmarks
We selected benchmarks that are close to real use cases, from a varied range of categories: reading comprehension, common sense, linguistic proficiency, knowledge, maths & reasoning, and bias. For each benchmark, where possible, we evaluated the Basque, English, and Spanish versions to facilitate the analysis of language-specific tradeoffs for each fine-tuned model variant. We included Spanish due to its close sociolinguistic relationship with Basque and to assess how our instruction-tuning methods affect a third language not directly targeted by our experiments. Thus, we evaluated a total of 29 benchmarks, as detailed in Section C.1.
For conducting these evaluations, we relied on Eleuther AI’s LM Evaluation Harness Biderman et al. (2024). Most datasets are framed as multiple-choice problems where models’ answers are determined by selecting the option with the highest log probability. For generative tasks, answers are directly sampled from the model. To provide models with contextual examples, our evaluations employed a few-shot setting. All results are measured for accuracy following standard, public implementations.999We implemented some benchmarks that are not currently available, and which will be contributed to the main repository (github.com/EleutherAI/lm-evaluation-harness) upon publication. Refer to Section C.1 for details.
@@ 5.2 Human Evaluation: Arena
Unlike static benchmarks, which rely on fixed datasets and automatic metrics, arena-style evaluations are better suited for assessing open-ended text generation, where subjective quality judgments play a central role. In this section, we first describe our implementation of the arena framework, including details on participants and evaluation conditions. For additional details about the human evaluation, refer to Section C.2, where we describe our infrastructure and introduce the Bradley–Terry model Bradley and Terry (1952), which we use to infer a model ranking from the collected pairwise preferences.
To gather human preferences for our evaluation, we organized a community-driven initiative. This collaborative effort ran for 14 days and attracted approximately 1,680 participants, resulting in a total of 12,890 preference annotations. The event was open to any Basque speaker, regardless of their proficiency level. Participants were required to register their educational background and language proficiency before contributing. Once registered, users could submit prompts and compare model responses. Fig. 2 shows that the majority of the participants—and, therefore, the preferences—have a bachelor or superior education and a high or native language proficiency level.
@@ 6 Results
In the following paragraphs, we disclose the effect of each component: backbone models, Basque corpora and bilingual instructions. For both the benchmark and the human evaluation, main results are presented in Fig. 3. Complete results for all tested models, broken down by benchmark and language, along with detailed arena evaluation results, can be consulted in Appendix D.
@@ 6.1 The Impact of Basque Corpora
We begin by analyzing the influence of Corpus EU, as it is intuitively the most critical resource for teaching a new language to an LLM. Fig. 3 confirms that this intuition aligns well with empirical results. In both human evaluations (Fig. 3(a)) and benchmark scores (Fig. 3(b)), models trained on Corpus EU achieved significantly better performance. The advantage is especially pronounced—up to 12 points in accuracy and over 300 points in arena score—when no other Basque signal (e.g., Basque instructions) is included. However, for models that already use Base EU as their backbone, additional exposure to Corpus EU offered little benefit and was sometimes even detrimental.
@@ 6.2 The Impact of Instruction Data
We analyze the effect of instruction data by comparing variants trained with no instructions, English-only instructions, Basque-only instructions, and a combination of both.
Starting with the question of whether to use instructions at all, we focus on Instruct EN-based variants. Human evaluation results (Fig. 3(a)) clearly show that incorporating instructions, regardless of language, helps mitigate catastrophic forgetting and improves arena scores by nearly 100 points. In contrast, benchmark results (Fig. 3(b)) show only marginal gains, particularly when using Basque instructions. This discrepancy highlights the limitations of static benchmarks and underscores the value of human or text generation-based evaluation.
When comparing instruction languages, we observe a general trend: English instructions tend to yield better results across both evaluation methods. However, there are exceptions. For example, models based on Instruct EN perform comparably or slightly better with Basque instructions in human evaluations, while Base EU-based models perform similarly on benchmarks. Notably, combining English and Basque instructions consistently produces the best results across most scenarios.
@@ 6.3 The Impact of Backbone Models and Curriculum Learning
By analyzing models trained from different backbones, we explore various curriculum learning strategies: (i) teaching the language first and then instruction-following, (ii) teaching instruction-following in English first and then the target language, or (iii) learning everything simultaneously.
@@
When comparing models based on Base EN (learning everything together) with those based on Base EU (learning the language first), we observe no significant difference in performance. Interestingly, the Base EU variant without access to Corpus EU achieves performance nearly identical to the Base EN variant with access to Corpus EU, across all instruction settings. This suggests that teaching the language in a separate pre-training step offers no measurable advantage. From this, we conclude that there is no compelling reason to separate language acquisition from instruction tuning.
@@
While instruction tuning pipelines are often complex and multi-staged Lambert et al. (2025), our approach adopts a simpler structure. Previous work by Xu et al. (2025) showed that models initialized from Base EN and trained with sampled instructions from Instruct EN perform comparably to Instruct EN. However, our findings indicate that this strategy does not transfer well to low-resource languages. As shown in Fig. 3, models based on Instruct EN consistently outperform those based on Base EN, both in human evaluations and benchmarks. These results support the conclusion that starting from a well-instructed English backbone yields better performance than learning everything from scratch.
@@ 6.4 The Impact of the Scale
Based on our previous analyses, we scaled up the variant that performed best in preliminary benchmark evaluations.101010We did this before running the arena as we wanted to include a 70B model in the human evaluation. Table 1 shows the results for some multiple-choice benchmarks and arena scores. On the one hand, we have the three best-performing 8B variants and the baseline. On the other hand, we present the results for the 70B best-performing variant and the baseline. Finally, we also show the performance of Claude 3.5 Sonnet and GPT-4o.
@@ 7 Analysis and Discussion
We find a strong correlation between Basque benchmark performance and human preferences in the arena, with Spearman’s rank correlation coefficients exceeding 0.80 across all arena dimensions (see Table 2), which suggests that automated Basque benchmarks may provide a reliable proxy for human evaluation in future research. English benchmarks display positive but non-significant correlations, while Spanish benchmarks show virtually no correlation with our Basque arena results. It must be noted that these correlations are not directly comparable across languages due to differences in benchmark composition. Detailed correlations by specific dataset can be found in Section D.1.
width=colspec=lX[2,r]X[3,l]X[2,r]X[3,l]X[2,r]X[3,l],
cells=font=,
row1=font=,
column1=font=,
column1=rightsep=3pt,
column2,4,6=rightsep=2pt,
column3,5,7=leftsep=1pt, rightsep=3pt,
rowsep=0pt
[…]
Bench EN 0.25 P = 0.32 0.23 P = 0.37 0.27 P = 0.29
Bench ES -0.02 P = 0.95 -0.09 P = 0.72 -0.04 P = 0.89
[…]
@@
Fig. 4 shows performance changes across languages relative to each backbone model on the multilingual benchmarks Belebele, MGSM, MMLU, and XStoryCloze. We observe that Instruct EN-based models exhibit a clear trade-off: improvements in Basque come at a cost of decreased performance in English and Spanish, suggesting a competitive relationship between languages in the model’s parameter space. In contrast, greater flexibility for multilingual adaptation is observed in Base EN models, which improve across all three languages (though from a lower absolute performance baseline). Base EU models show moderate changes with gains primarily in Spanish and English rather than Basque. As observed previously, Corpus EU consistently yields the largest performance gains for the target language. Among Instruct EN variants, the configuration with Corpus EU and Instructions EN achieves the most Basque improvement and the least regression in other languages. Despite these adaptation strategies, models still perform better in English and Spanish on equivalent tasks, with the exception of culturally-specific knowledge as evidenced by the results on the BertaQA dataset (see complete results in Section D.2).
@@
This paper does not address the alignment step typically required to reduce biases and prevent unsafe responses in production-ready LLMs. However, the backbone model Instruct EN is already aligned, and we expect variants based on this backbone to retain safety-related behaviors. To verify this, we conducted a small-scale experiment to assess the safety and bias characteristics of our models after the language adaptation (see Appendix E). Results indicate that the models preserve most of the safety and bias alignment properties of Instruct EN in English, and that these properties effectively transfer to Basque, with no observed degradation after adaptation.

### v2

@@ Abstract
Instructing language models with user intent requires large instruction datasets, which are only available for a limited set of languages. In this paper, we explore alternatives to conventional instruction adaptation pipelines in low-resource scenarios. We assume a realistic scenario for low-resource languages, where only the following are available: corpora in the target language, existing open-weight multilingual base and instructed backbone LLMs, and synthetically generated instructions sampled from the instructed backbone. We present a comprehensive set of experiments for Basque that systematically study different combinations of these components evaluated on benchmarks and human preferences from participants. Our conclusions show that target language corpora are essential, with synthetic instructions yielding robust models, and, most importantly, that using as backbone an instruction-tuned model outperforms using a base non-instructed model. Scaling up to Llama 3.1 Instruct 70B as backbone, our model comes near frontier models of much larger sizes for Basque, without using any Basque instructions. We release code, models, instruction datasets, and human preferences to support full reproducibility in future research on low-resource language adaptation.111github.com/hitz-zentroa/latxa-instruct
@@ 1 Introduction
Specifically, this work systematically explores diverse strategies beyond the conventional pipeline for developing instruction-tuned models for low-resource languages, seeking to identify optimal adaptation paths for Basque as our primary case study (see Fig.˜1). We deliberately constrain our exploration to resources either readily available or creatable using open models, avoiding reliance on distillation from commercial state-of-the-art systems. While our investigation focuses on a single language, our findings likely generalize to many similarly-resourced languages: Basque represents an ideal test case, ranking approximately 50th in Common Crawl with a presence approximately times smaller than English,222commoncrawl.github.io/cc-crawl-statistics/plots/languages.html and notably lacking pre-existing instruction datasets. This resource profile mirrors challenges faced by numerous other low-resource languages worldwide.
@@ 2 Related Work
Research on developing LLMs for under-resourced languages has explored various approaches, with varying degrees of success. Initial attempts to develop models from scratch for specific low-resource languages have proven challenging due to limited training data. Multilingual model development has emerged as a more promising strategy, with researchers leveraging cross-lingual transfer learning to improve performance Scao et al. (2023); Le Scao et al. (2022); Shliazhko et al. (2024). The most effective approach to date involves continued pretraining of existing multilingual models, which allows for language-specific adaptation while benefiting from the rich linguistic representations of larger training corpora Etxaniz et al. (2024b); Luukkonen et al. (2023); Tran et al. (2024). While progress has been made in developing these base models, optimal methods for instructing and fine-tuning them for under-resourced languages remain largely unexplored Gonzalez-Agirre et al. (2025); Martins et al. (2025); Üstün et al. (2024).
@@ 3 Resources
Instructing LLMs typically relies on two components: base (or foundational) LLMs and instruction datasets. For non-hegemonic languages, obtaining instruction datasets can be very challenging, particularly in low-resource language scenarios. In the case of Basque in particular, there are no manually generated, or even good quality automatically generated, large sets of instruction-answer pairs. Consequently, as shown in Fig.˜1, our available resources are constrained to corpora on the target language, and base and instruct models for high-resource languages. From these limited resources, we derive the necessary components to create Basque instruction-tuned models through strategic combinations of synthetic data generation and model adaptation. In the following sections, we describe these seed resources and derivations.
@@ 3.1 Basque corpora
For the pretraining data, we have leveraged the corpora used to train Latxa, the first family of LLMs trained specifically for Basque Etxaniz et al. (2024b). This corpus comprises 4.3M of high-quality documents in Basque, roughly 3.5B Llama 3.1 tokens. Among the sources, it contains high-quality news data extracted using ad-hoc scrapers Artetxe et al. (2022), Wikipedia333The 20231101 dump corresponding to November 2023. and sources based on Common Crawl such as CulturaX Nguyen et al. (2024), Colossal OSCAR Abadji et al. (2022) and HLPT v1.1 de Gibert et al. (2024). This corpus comes normalized, deduplicated and filtered. The data is publicly available in the HuggingFace hub.444hf.co/datasets/HiTZ/latxa-corpus-v1.1 We will henceforth refer to this corpus as Corpus EU.
@@ 3.3 Instruction Sampling and Translation
To avoid external influences, we instead sample instructions directly from our InstructEN model. We generate the English instructions following Xu et al. (2025). Using this technique, we conditioned InstructEN to generate instructions of different types and tasks: general-purpose, code, math, arithmetic and translation. We generated a total of 4M English instructions. However, after a hyperparameter search, we found out that using just 1M instructions yielded better results overall (see Appendix˜B). We share more details and examples of the process in Appendix˜A.
@@
We translated instructions sampled from InstructEN using few-shot prompting with Base EU. Existing machine translation systems for the English–Basque language pair (e.g., NLLB Costa-jussà et al. (2022)) are primarily trained on sentence-level textual data and often struggle with more complex inputs, including selectively translating natural language content embedded within code snippets. By leveraging an LLM like Base EU, which has been exposed to diverse data types, we obtained higher-quality translations for this setting. Moreover, using a model trained within our own experimental framework allows us to avoid introducing external factors into our pipeline. More details about the process and prompts used to translate the instructions are given in Appendix˜A.
@@ 4 Experimental Setup
We formalize our experimental setup as follows. Let be the set of backbone models and be the set of binary variables indicating whether to use Basque corpora, English instructions, and/or Basque instructions. The space of possible configurations is thus , where is the power set of , yielding theoretical combinations. Note that we explore training strategies that leverage both raw text and instruction data simultaneously. From the total of 24 combinations, we exclude redundant configurations where a model is retrained on data it was originally trained with. The resulting set of distinct instruction-tuned model variants therefore comprises 18 configurations: the original Llama 3.1 Instruct 8B (i.e., Instruct EN) and 17 new 8B-sized models. Table˜3 in Appendix˜B provides a complete account of all model variants and their shorthand names. Additionally, we trained a 70B model following the configuration that performed best in preliminary benchmark evaluations.
@@ 5 Evaluation
We employed two complementary evaluation approaches to assess the impact of each instruction-tuning recipe. On the one hand, we used a selection of static benchmarks that evaluate specific model capabilities and knowledge through standardized tests. On the other hand, we conducted human evaluations through A/B testing (arena style) to capture qualitative aspects of model performance. In addition, we look into the impact of our recipes on safety and bias.
@@ 5.1 Static Benchmarks
We selected benchmarks that are close to real use cases, from a varied range of categories: reading comprehension, common sense, linguistic proficiency, knowledge and maths & reasoning.
For each benchmark, where possible, we evaluated the Basque, English, and Spanish versions to facilitate the analysis of language-specific tradeoffs for each fine-tuned model variant. This choice of evaluation languages reflects the linguistic reality of the Basque-speaking community in northern Spain, where Basque and Spanish are co-official and English is the most commonly taught foreign language. Importantly, these languages come from distinct families: Basque is a language isolate, Spanish is Romance, and English is Germanic. Thus, we examine cross-lingual transfer effects and assess whether improvements in our language come at the cost of performance in related community languages, including one—Spanish—not directly targeted by our experiments. In total, then, we evaluated 27 benchmarks, as detailed in Section˜C.1.
For conducting these evaluations, we relied on LM Evaluation Harness Biderman et al. (2024). Most datasets are framed as multiple-choice problems where models’ answers are determined by selecting the option with the highest log probability. For generative tasks, answers are directly sampled from the model. To provide models with contextual examples, our evaluations employed a few-shot setting. All results are measured for accuracy following standard, public implementations. Refer to Section˜C.1 and our repository for details.
@@ 5.2 Human Evaluation: Arena
Unlike static benchmarks, which rely on fixed datasets and automatic metrics, arena-style evaluations are better suited for assessing open-ended text generation, where subjective quality judgments play a central role. In this section, we first describe our implementation of the arena framework, including details on participants and evaluation conditions. For additional details about the human evaluation, refer to Section˜C.2, where we describe our infrastructure and introduce the Bradley–Terry model Bradley and Terry (1952), which we use to infer a model ranking from the collected pairwise preferences.
To gather human preferences for our evaluation, we organized a community-driven initiative. This collaborative effort ran for 14 days and attracted approximately 1,680 participants, resulting in a total of 12,890 preference annotations. The event was open to any Basque speaker, regardless of their proficiency level. Participants were required to register their educational background and language proficiency before contributing. Once registered, users could submit prompts and compare model responses. Fig.˜2 shows that the majority of the participants—and, therefore, the preferences—have a bachelor or superior education and a high or native language proficiency level.
@@
5.3 Safety and Bias
This paper does not address the alignment step typically required to reduce biases and prevent unsafe responses in production-ready LLMs. However, the backbone model Instruct EN is already aligned, and we expect variants based on this backbone to retain safety-related behaviors. To verify this, we constructed a new Basque-English parallel dataset inspired by XSTest Röttger et al. (2024), combining clearly unsafe prompts with superficially similar but safe ones. We measure both Violation rates (VR) and False Refusal rates (FRR), where the model wrongly declines safe prompts. For bias, we rely on BBQ Parrish et al. (2022) and its adaptation to Basque Zulaika and Saralegi (2025), reporting results in terms of accuracy. Further details are available in Section˜C.3.
@@ 6 Results
In the following paragraphs, we disclose the effect of each component: backbone models, Basque corpora and bilingual instructions. For both the benchmark and the human evaluation, main results are presented in Fig.˜3. Complete results for all tested models, broken down by benchmark and language, along with detailed arena evaluation results, can be consulted in Appendix˜D.
@@ 6.1 The Impact of Basque Corpora
We begin by analyzing the influence of Corpus EU, as it is intuitively the most critical resource for teaching a new language to an LLM. Fig.˜3 confirms that this intuition aligns well with empirical results. In both human evaluations (Fig.˜3(a)) and benchmark scores (Fig.˜3(b)), models trained on Corpus EU achieved significantly better performance. The advantage is especially pronounced—up to 12 points in accuracy and over 300 points in arena score—when no other Basque signal (i.e., Basque instructions) is included. However, for models that already use Base EU as their backbone, additional exposure to Corpus EU offered little benefit and was sometimes even detrimental.
@@ 6.2 The Impact of Instruction Data
We analyze the effect of instruction data by comparing variants trained with no instructions (), English-only instructions (I EU), Basque-only instructions (I EN), and their combination (I EN+EU).
Starting with the question of whether to use instructions at all, we focus on Instruct EN-based variants. Human evaluation results (Fig.˜3(a)) clearly show that incorporating instructions, regardless of language, helps mitigate catastrophic forgetting and improves arena scores by nearly 100 points. In contrast, benchmark results (Fig.˜3(b)) show only marginal gains, particularly when using Basque instructions. This discrepancy highlights the limitations of static benchmarks and underscores the value of human or text generation-based evaluation.
When comparing instruction languages, we observe a general trend: English instructions tend to yield better results across both evaluation methods. However, there are exceptions. For example, models based on Instruct EN perform comparably or slightly better with Basque instructions in human evaluations, while Base EU-based models perform similarly on benchmarks.
Notably, combining English and Basque instructions consistently produces the best results across most scenarios. While this improvement could be attributed to certain model variants having access to more training data, our preliminary results in Appendix˜B refute this hypothesis, as using more monolingual instructions (1M vs 4M) resulted in similar results.
@@ 6.3 The Impact of Backbone Models and Curriculum Learning
By analyzing models trained from different backbones, we explore various curriculum learning strategies: (i) teaching the language first and then instruction following, (ii) teaching instruction following in English first and then the target language, or (iii) learning everything simultaneously.
@@
When comparing models based on Base EN (i.e., acquiring Basque and instruction-following capabilities simultaneously) with those based on Base EU (learning the language first, then instruction following), we observe no significant difference in performance. Interestingly, the Base EU variant without access to Corpus EU during instruction tuning achieves performance nearly identical to the Base EN variant with access to Corpus EU, across all instruction settings. This suggests that teaching the language in a separate pretraining step offers no measurable advantage. From this, we conclude that there is no compelling reason to separate language acquisition from instruction tuning.
@@
While instruction tuning pipelines are often complex and multi-staged Lambert et al. (2025), our approach adopts a simpler structure. Previous work by Xu et al. (2025) showed that models initialized from Base EN and trained with sampled instructions from Instruct EN perform comparably to Instruct EN. However, our findings indicate that this strategy does not transfer well to low-resource languages. As shown in Fig.˜3, models based on Instruct EN consistently outperform those based on Base EN, both in human evaluations and benchmarks. These results support the conclusion that starting from a well-instructed English backbone yields better performance than learning everything from scratch.
@@ 6.4 The Impact of the Scale
Based on our previous analyses, we scaled up the variant that performed best in preliminary benchmark evaluations.999We did this before running the arena as we wanted to include a 70B model in the human evaluation. Table˜1 shows the results for some multiple-choice benchmarks and arena scores. On the one hand, we have the three best-performing 8B variants and the baseline. On the other hand, we present the results for the 70B best-performing variant and the baseline. Finally, we also show the performance of Claude 3.5 Sonnet and GPT-4o.
@@ 7 Analysis and Discussion
Fig.˜4 shows Spearman’s rank correlation coefficients, , between benchmark performance and arena scores, across different benchmark languages (Basque, English, Spanish) and arena dimensions (content, language, and global). We observe that average Basque benchmark performance and arena rankings correlate strongly, with across all arena dimensions—which suggests that automated Basque benchmarks may provide a reliable proxy for human evaluation in future research. This correlation is particularly pronounced for specific benchmarks, including EusProficiency, EusTrivia, EusExams and BertaQA, which are interestingly the datasets that were natively constructed in Basque, rather than translated from existing English benchmarks. The average of English benchmarks displays overall positive but non-significant correlations. Only BertaQA shows positive correlations, with the local subset obtaining correlations similar to the Basque BertaQA, likely reflecting the types of culturally-specific questions that users posed in the arena. Spanish benchmarks show, on average, no correlation with the arena. Some of the English and Spanish benchmarks show large negative correlations, reflecting the performance trade-off between Basque and other languages.
@@
Fig.˜5 shows performance changes across languages relative to each backbone model on the multilingual benchmarks Belebele, MGSM, MMLU, and XStoryCloze. We observe that Instruct EN-based models exhibit a clear trade-off: improvements in Basque come at a cost of decreased performance in English and Spanish, suggesting a competitive relationship between languages in the model’s parameter space. In contrast, greater flexibility for multilingual adaptation is observed in Base EN models, which improve across all three languages (though from a lower absolute performance baseline). Base EU models show moderate changes with gains primarily in Spanish and English rather than Basque. As observed previously, Corpus EU consistently yields the largest performance gains for the target language. Among Instruct EN variants, the configuration with Corpus EU and Instructions EN achieves the most Basque improvement and the least regression in other languages. Despite these adaptation strategies, models still perform better in English and Spanish on equivalent tasks, with the exception of culturally-specific knowledge as evidenced by the results on the BertaQA dataset (see complete results in Section˜D.1).
@@
We evaluate the model variant INS EN C EU I EN, comparing it with two critical counterparts: (1) BAS EN C EU I EN, to analyze the impact of starting from an already instruction-tuned backbone versus a base model (8B only); and (2) Instruct EN, to measure potential alignment changes introduced with our fine-tuning data mix (8B and 70B). As shown in Table˜2, the BAS EN C EU I EN model demonstrates low safety and high bias in the Basque language compared to English, reflecting a significant lack of alignment. In contrast, Instruct EN and its variant INS EN C EU I EN are better aligned in both languages. Interestingly, in Basque, the 8B Instruct EN shows a lower VR than both BAS EN C EU I EN and INS EN C EU I EN models and even the larger 70B counterpart. This behavior is due to the model’s limited comprehension of instructions in the Basque language, leading it to produce inconsistent yet safe responses. On the other hand, the 70B models display comparable performance across both languages, with the INS EN C EU I EN version slightly outperforming others and achieving a lower VR in Basque. Notably, the FRR remains very low across all models, indicating that safety mechanisms do not come at the cost of excessive conservatism. Bias outcomes in English are consistently better than in Basque, and larger models generally perform better than smaller ones. However, the differences in bias between the original and language-adapted models are minimal. Overall, we demonstrate that much of the safety and bias alignment is transferred from Instruct EN, not only in the newly added language, but also in the predominant language.
| EU | EN | ||||||
| VR | FRR | BBQ | VR | FRR | BBQ | ||
| 8B | INS EN C EU I EN | 24.00 | 0.00 | 70.80 | 16.00 | 0.00 | 87.06 |
| BAS EN C EU I EN | 44.00 | 0.00 | 51.73 | 20.00 | 4.00 | 72.35 | |
| Instruct EN | 8.00 | 0.00 | 71.03 | 4.00 | 0.00 | 87.65 | |
| 70B | INS EN C EU I EN | 4.00 | 0.00 | 85.29 | 4.00 | 0.00 | 94.38 |
| Instruct EN | 20.00 | 0.00 | 84.90 | 8.00 | 0.00 | 95.32 |